# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ 3.2 Global convergence
Firstly, from Proposition 3.7 we can conclude that there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that the sequence $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ follows the update scheme
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}$ $=\frac{1}{d}{\bm{Z}}_{k}{\bm{W}}{\bm{1}}_{d}-\eta_{k}\frac{1}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}}_{d}$ (3.10)
$\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}\left(\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right).$
Moreover, Assumption 3.1(2) illustrates that
$\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda_{\eta}(\lambda_{\eta}(i)+T)}\left\|\sum_{k=i}^{j}\frac{\eta_{k}}{d}\Xi_{k+1}{\bm{1}}_{d}\right\|$ (3.11)
[…]
Then combined with Assumption 3.1(1), it directly follows from [36, Lemma 2.20] that the interpolated process of the sequence $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution of the differential inclusion (3.2). This completes the proof.
Based on Proposition 3.8 and Theorem 2.10, we establish the convergence theorem, where we prove that $\{{\bm{z}}_{i,k}\}$ converges subsequentially to the stable set of (3.2) for our framework (DSM).
Theorem 3.9.
Suppose Assumption 3.1 and Assumption 3.3 hold. Then for the sequence $\{{\bm{Z}}_{k}\}$ generated by our framework (DSM), any cluster point of $\{{\bm{Z}}_{k}\}$ lies in $\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{z}}{\bm{1}}_{d}^{\top},~{}{\bm{z}}\in\mathcal{A}\}$ and the sequence $\{\psi({\bm{z}}_{i,k})\}$ converges for each $i\in[d]$.
Proof.
Proposition 3.8 illustrates that the interpolated process of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution to the differential inclusion (3.2). Together with Assumption 3.3 and Theorem 2.10, we can conclude that any cluster point of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ lies in $\mathcal{A}$, and the sequence of function values $\{\psi(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})\}$ converges.
Moreover, notice that Proposition 3.6 illustrates that $\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{P}}\right\|=0$. Therefore, we can conclude that any cluster point of $\{{\bm{Z}}_{k}\}$ lies in $\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{z}}{\bm{1}}_{d}^{\top},~{}{\bm{z}}\in\mathcal{A}\}$, and the sequence of function values $\{\psi({\bm{z}}_{i,k})\}$ converges for each $i\in[d]$.
4 Application: Decentralized Stochastic Subgradient-based Methods
In this section, we aim to illustrate the great flexibility of our proposed framework (DSM). We show that our proposed framework (DSM) encloses a wide range of popular decentralized stochastic subgradient-based methods, including DSGD, DSGD with momentum (DSGD-M), and DSGD with gradient-tracking technique (DSGD-T). More importantly, we demonstrate that our theoretical results in Theorem 3.9 provide theoretical convergence guarantees for these decentralized stochastic subgradient-based methods, especially in the minimization of nonsmooth path-differentiable functions.
Throughout this section, we make the following assumptions on (DOP).
@@ 4 Application: Decentralized Stochastic Subgradient-based Methods
    For each $i\in[d]$ and ${\bm{s}}_{l}\in\mathcal{S}_{i}$, $F(\cdot;{\bm{s}}_{l})$ is definable and admits a definable conservative field $D_{F(\cdot;{\bm{s}}_{l})}$.

### v3

@@ 3.1 Basic assumptions and main results
From Theorem 3.5, there exists $r$, $\alpha_{\text{ub}},T>0$, for any $\alpha_{\text{ub}}$-upper-bounded $\{\eta_{k}\}$ and $\{\delta_{k}\}$, and any $(\alpha_{ub},T,\{\eta_{k}\})$-controlled sequence $\{\Xi_{k+1}\}$, the sequence $\{{\bm{Z}}_{k}\}\subseteq\mathcal{L}_{2r}^{d}$. Taking sufficiently large $M$, (3.2) becomes
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}[\Phi^{s_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}],$
where $s_{k}:=\delta_{k}+\|{\bm{Z}}_{\perp,k}\|$. Since $\lim_{k\to\infty}\|{\bm{Z}}_{\perp,k}\|\leq\lim_{k\to\infty}\sum_{i=0}^{k-1}\lambda_{2}^{k-1-i}\|\eta_{i}({\bm{H}}_{i}+\Xi_{i+1})\|=0$, we have $\lim_{k\to\infty}s_{k}=0$.
Moreover, $(\alpha_{ub},0,T,\{\eta_{k}\})$-asymptotically controlled sequence $\{\Xi_{k+1}\}$ implies that for any $T^{\prime}>0$,
$\lim_{s\to\infty}\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T^{\prime})}\left\|\sum_{k=s}^{i}\eta_{k}\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right\|=0,$
then it follows that the interpolated process of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution of (3.1).
By combining Lemma 2.16, we can infer that, any cluster point of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ lies in $\mathcal{A}$ and the sequence $\{\psi(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})\}$ converges. As $\lim_{k\to\infty}\|{\bm{Z}}_{\perp,k}\|=0$, we conclude that, any cluster point of $\{{\bm{z}}_{i,k}\}$ lies in $\mathcal{A}$ and $\{\psi({\bm{z}}_{i,k})\}$ converges, for any $i\in[d]$. This completes the proof. ∎
3.2 Convergence with random reshuffling
Given a family of locally bounded, graph-closed set-valued mappings $\{\mathcal{U}_{i,j}:1\leq i\leq d,0\leq j\leq N-1\}$, we consider framework (DSM) with the following random reshuffled update scheme,
$\frac{1}{d}\Xi_{k+1}\mathbf{1}_{d}\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{U}_{i,i_{k}}^{\rho_{k}}({\bm{z}}_{i,k})-\frac{1}{d}{\bm{H}}_{k}\mathbf{1}_{d}.$ (3.8)
One can see (3.8) as a detailed characterization of the relationship between the update direction and the iterates $\{{\bm{z}}_{i,k}\}$, thereby serving as a replacement for the role of (1.5). Hereafter, we make some assumptions on (3.8) and step-sizes $\{\eta_{k}\}$ in (DSM).
@@ 3.2 Convergence with random reshuffling
    For any $k_{1},k_{2}\in\mathbb{N}$, $\eta_{k_{1}}=\eta_{k_{2}}$ holds, if $\lfloor\frac{k_{1}}{N}\rfloor=\lfloor\frac{k_{2}}{N}\rfloor$.